# Bounded Rationality Equilibrium Learning in Mean Field Games

### v1

@@ Acknowledgements
This work has been co-funded by the LOEWE emergenCITY research promotion program of the federal state of Hessen, Germany, by the German Research Foundation (DFG) within the Collaborative Research Center (CRC) 1053 MAKI, by the Federal Ministry of Education and Research as part of the Software Campus project RL4MFRP (funding code 01IS23067) and by the Hessian Ministry of Science and the Arts (HMWK) within the projects ”The Third Wave of Artificial Intelligence - 3AI” and hessian.AI.
Appendix A Proof of Proposition 2
@@
Appendix B Proof of Proposition 3
@@
For BE and RE, the result is known from (Cui & Koeppl, 2021, Thm. 3). For QRE, the proof follows similarly by obtaining Lipschitz continuity of the fixed point map. In particular, the map $\Gamma_{\mathcal{M}}$ from policy to MF is Lipschitz by Assm. 2. Similarly, the map $\pi\mapsto\Gamma_{Q}(\Gamma_{\mathcal{M}}(\pi),\pi)$ mapping to the state-action value function is Lipschitz. Lastly, the map $\Gamma^{\alpha}_{\Pi}$ is Lipschitz with a constant that is inversely proportional to $\alpha$ (Gao & Pavel (2017, Prop. 4) or Cui & Koeppl (2021, Lemma B.7.5)). Therefore, for sufficiently large $\alpha>0$, we have that $\pi\mapsto\Gamma^{\alpha}_{\Pi}(\Gamma_{Q}(\Gamma_{\mathcal{M}}(\pi),\pi))$ is a contraction. ∎
Appendix C Proof of Theorem 2
@@
where $\mu^{\pi^{*,\tau^{\prime}}}=\Gamma_{\mathcal{M}}(\pi^{*,\tau^{\prime}})$ is generated by the best regularized response against the current MF $\bar{\mu}^{\tau^{\prime}}$, i.e. $\pi^{*,\tau^{\prime}}=\Gamma^{\alpha}_{\Pi}(\Gamma^{\alpha}_{\tilde{Q}}(\bar{\mu}^{\tau^{\prime}})$ for all $\tau^{\prime}>1$. The initial values for $\tau\leq 1$ are given by some arbitrary initial policy and its MF, as in (Perrin et al., 2020).
@@
where the latter equality follows by dropping derivatives w.r.t. $\pi^{*,\tau}$ via envelope theorem (Afriat, 1971), since $\mathcal{E}^{\mathrm{RE}}(\bar{\pi}^{\tau})$ is continuously differentiable in $\tau$ by continuity of the maps $\Gamma^{\alpha}_{\Pi}$, $\Gamma^{\alpha}_{\tilde{Q}}$ and hence also (8).
@@
$\quad=\alpha\sum_{t\in\mathcal{T}}\sum_{x\in\mathcal{X}}\frac{1}{\tau}\mu^{\pi^{*,\tau^{\prime}}}_{t}(x)\left(\sum_{u\in\mathcal{U}}\pi^{*,\tau}_{t}(u\mid x)-\sum_{u\in\mathcal{U}}\bar{\pi}^{\tau}_{t}(u\mid x)\right)=0.$
Moreover, for the first and fourth term, w.l.o.g. let $r$ not depend on actions, then we have
@@
Appendix D Definitions and Notations for RHE
@@
Appendix E Proof of Theorem 3
@@ G.3 SIS Game
Appendix H Additional Experiments

### v2

@@ Acknowledgements
This work has been co-funded by the LOEWE emergenCITY research promotion program of the federal state of Hessen, Germany, by the German Research Foundation (DFG) within the Collaborative Research Center (CRC) 1053 MAKI and project number 517777863, by the Federal Ministry of Education and Research as part of the Software Campus project RL4MFRP (funding code 01IS23067) and by the Hessian Ministry of Science and the Arts (HMWK) within the projects ”The Third Wave of Artificial Intelligence - 3AI” and hessian.AI.
Appendix A A  Proof of Proposition 2
@@
Appendix B B  Proof of Proposition 3
@@
For BE and RE, the result is known from (Cui and Koeppl 2021, Thm. 3). For QRE, the proof follows similarly by obtaining Lipschitz continuity of the fixed point map. In particular, the map $\Gamma_{\mathcal{M}}$ from policy to MF is Lipschitz by Assm. 2. Similarly, the map $\pi\mapsto\Gamma_{Q}(\Gamma_{\mathcal{M}}(\pi),\pi)$ mapping to the state-action value function is Lipschitz. Lastly, the map $\Gamma^{\alpha}_{\Pi}$ is Lipschitz with a constant that is inversely proportional to $\alpha$ (Gao and Pavel (2017, Prop. 4) or Cui and Koeppl (2021, Lemma B.7.5)). Therefore, for sufficiently large $\alpha>0$, we have that $\pi\mapsto\Gamma^{\alpha}_{\Pi}(\Gamma_{Q}(\Gamma_{\mathcal{M}}(\pi),\pi))$ is a contraction. ∎
Appendix C C  Proof of Theorem 2
@@
where $\mu^{\pi^{*,\tau^{\prime}}}=\Gamma_{\mathcal{M}}(\pi^{*,\tau^{\prime}})$ is generated by the best regularized response against the current MF $\bar{\mu}^{\tau^{\prime}}$, i.e. $\pi^{*,\tau^{\prime}}=\Gamma^{\alpha}_{\Pi}(\Gamma^{\alpha}_{\tilde{Q}}(\bar{\mu}^{\tau^{\prime}})$ for all $\tau^{\prime}>1$. The initial values for $\tau\leq 1$ are given by some arbitrary initial policy and its MF, as in (Perrin et al. 2020).
@@
where the latter equality follows by dropping derivatives w.r.t. $\pi^{*,\tau}$ via envelope theorem (Afriat 1971), since $\mathcal{E}^{\mathrm{RE}}(\bar{\pi}^{\tau})$ is continuously differentiable in $\tau$ by continuity of the maps $\Gamma^{\alpha}_{\Pi}$, $\Gamma^{\alpha}_{\tilde{Q}}$ and hence also (8).
@@
$\quad=\alpha\sum_{t\in\mathcal{T}}\sum_{x\in\mathcal{X}}\frac{1}{\tau}\mu^{\pi^{*,\tau^{\prime}}}_{t}(x)\left(\sum_{u\in\mathcal{U}}\pi^{*,\tau}_{t}(u\mid x)-\sum_{u\in\mathcal{U}}\bar{\pi}^{\tau}_{t}(u\mid x)\right)=0,$
where the last equation follows from the fact that the sum of the probabilities assigned by a policy to all possible actions is equal to one. Moreover, for the first and fourth term, w.l.o.g. let $r$ not depend on actions, then we have
@@
Appendix D D  Definitions and Notations for RHE
@@
Appendix E E  Proof of Theorem 3
@@ Appendix G G  Benchmarks
Appendix H H  Additional Experiments